OpenStudy (anonymous):

Solve the proportion x over x squared minus 2 equals 2 over 3 times x plus 4.. What is the value of x?

OpenStudy (anonymous):

\[{x \over x^2} {- 2} ={2 \over 3} \times {x+4}\] Is this the question?

OpenStudy (anonymous):

\[\frac{ x }{ x ^{2}-2 }=\frac{ 2 }{ 3x+4}\]

OpenStudy (anonymous):

ahhh. Ok. Hint: Cross multiply.

OpenStudy (anonymous):

no i know that but i cant figure it out!!! its confusing and shiz

OpenStudy (anonymous):

3x^2+4x=2x^2-4 3x^2-2x^2+4x=-4 x^2+4x+4=0 Hint: Use quadratic formula to find x.

OpenStudy (anonymous):

yayayayyaayayayayay thats what i was looking for! i couldnt figure out how to get them all on one side! thankssssssssssssss :D

OpenStudy (anonymous):

you're welcome.